Solve an equation to find the length ofAB and BC.
B is between A and C. AB = x, BC = 2x, AC = 18

Answer by ikleyn(52803)

.
Solve an equation to find the length ofAB and BC.
B is between A and C. AB = x, BC = 2x, AC = 18
~~~~~~~~~~~~~~~~~~

Since B is between A and C, we have


    AB + BC = AC

     x + 2x = 18

      3x    = 18

       x    = 18/3 = 6.


ANSWER.  AB = 6 units;  BC = 2*6 = 12 units.

Solved.